P2-2 Transaction costs In late December you decide, for tax purposes, to sell a losing posi tion that you hold in Twitter, which is listed on the NYSE, so that you can capture the loss and use it to offset some capital gains, thus reducing your taxes for the current year However, since you still believe that Twitter is a good long-term investment, you wish to buy back your position in February the following year. To get this done you call your Charles Schwab brokerage account manager and request that he immediately sell your 1,200 shares of Twitter and then in early February buy them back. Charles Schwab charges a commission of $4.95 for online stock trades and for broker-assisted trades there is an additional $25.00 service charge, so the total commission is $29.95. a. Suppose that your total transaction costs for selling the 1,200 shares of Twitter in December were S59.95. What was the bid/ask spread for Twitter at the time b. Given that Twitter is listed on the NYSE, do your total transaction costs for c. When your February statement arrives in the mail, you see that your total trans- d. What are your total round-trip transaction costs for both selling and buying the your trade was executed? December seem reasonable? Explain why or why not. action costs for buying the 1,200 shares of Twitter were $47.95. What was the bid/ask spread for Twitter at the time your trade was executed? shares, and what could you have done differently to reduce the total costs?
Answer #1

a.

the bid/ask spread is difference between price the share bought(ask) and sold(bid) so,if sold is more than there is profit otherwise loss.So the transaction cost can be treated as spread for trade being executed.the bis/ask spread along with total brokerage /commission is defined as total transaction cost

so bid/ask spread=59.95-29.95

bid/ask spread=30 $

b.let us assume market price of buying is 10$ per share so cost of buying 1200 shares is 12000$

total commission paid is 29.95$.

so transaction cost =12029.95-12000

=29.95.which seems to be reasonable as per the calculations because it will also include spread.

c.transaction cost=47.95-29.95

=18$

d.Round up transaction cost=total cost of selling-total cost of buying

=59.95-47.95

=12.

so the total transaction cost will help to determine the total profit for the buying and selling.the investor will try to reduce the total cost by reducing other cost included in transaction cost like commission and service charges.